OpenStudy (anonymous):

find the values of c that make the following function continuous everywhere for 2x^2 + c if x<2 and cx^2+2 if x <2

OpenStudy (agent0smith):

The endpoints of the lines must meet - ie when x=2, \[\Large 2x^2 + c = cx^2+2\]

OpenStudy (anonymous):

so would you then write it out 8+c=c(4)+2? or where would you go once you know you have to set them equal?

OpenStudy (agent0smith):

That's right. Now find c from that.

OpenStudy (anonymous):

does it just come out to be c=2?

OpenStudy (agent0smith):

Yep. Which if you think about it, you can see it's 2 from here: \[\Large 2x^2 +c = cx^2 +2 \] from here, the 'x' parts must be equal, and so must the number parts... ie 2x^2 = cx^2 (ie c has to be 2) and c = 2